Olive Oil Olive oil contains triglycerides such as
Answer #1

Let us expand the long chains in the first compound (a).

CH2-O-CO-(CH2)14-CH3 this long chain of acid is saturated.

CH-O-CO-(CH2)7-CH=CH-(CH2)7-CH3 Contains one double bond, hence an unsaturated acid.

CH2-O-CO-(CH2)7-CH=CH-CH2-CH=CH-(CH2)4-CH3 Contains two double bonds hence unsaturated acid.

Compound (b)

CH2-O-CO-(CH2)7-CH=CH-(CH2)7-CH3 contains a double bond, hence unsaturated acid

CH-O-CO-(CH2)16-CH3 Saturated acid

CH2-O-CO-(CH2)7-CH=CH-(CH2)7-CH3 contains a double bond, hence unsaturated acid.

The unsaturated acids are likely to be in trans forms, where the bulky alkyl groups are far apart. If they are cis forms, the bulky alkyl groups will come closer, which causes the repulsion making this isomer less stable.